OpenStudy (anonymous):

Write the equation of the line given the slope and a point on the line. Type your answer in the box without skipping any spaces. m = 3, point (2, 1)

OpenStudy (anonymous):

The point slope form is given as y-y1=m(x-x1) where (x1,y1) is a point on the line. y-1 = 3 ( x-2) y= 3(x-2)+1

OpenStudy (goformit100):